A municipality wishes to determine the average utility bill for its county's population. In a random sample of 100 customers, it was found that the sample average utility bill was $112.59. Assuming the population standard deviation is $18.53, use Excel to find the margin of error of the confidence interval with a 97% confidence level.

Round your answer to two decimal places.
Answer #1

Solution:

We are given

n = 100

σ = 18.53

Confidence level = 97%

That is, c = 0.97, α = 1 - 0.97 = 0.03

α/2 = 0.03/2 = 0.015

1 - α/2 = 1 - 0.015 = 0.985

Margin of error = Z*σ/sqrt(n)

Excel formula: =normsinv(0.985)*18.53/sqrt(100)

Margin of error = 4.021177

Margin of error = 4.02